1. The cost of capital represents the weighted average cost of all sources of long-term financing to the firm, is normally the discount rate to use in analyzing an investment, is based on the valuation techniques from the previous chapter and is applied to bonds, preferred stock and common stock.What does this mean to you, in plain language? (2 paragraph well written clear explanation)

2. How do we measure the cost of preferred stock and cost of common stock equity of a firm? (1 concise paragraph 6 sentences minimum)
Answer #1

Fundamentally there are 4 sources of finance. Common stock, preferred stock, Debt and retained earnings. All these funds belong to external people, who have varied expectations as return for they parting with their money. Depending on class of investment they have different risk assumptions and expected returns.

Calculation of WACC implies, that we are trying to know the minimum percentage we have to earn if we have to justify the expectations of our investors. It depends on proportions of these funds in the capital structure and individual cost of capital.

WACC = Ke (%E) + Kd (%D) + Kp(%P) + Kr(%R)

Using the WACC as the discount rate for checking the financial feasibility of a project is therefore a welcome step. A project with positive NPV promises that the investors expectations are meet and there will be value addition to wealth of the common stock holder.

Investing without checking the WACC can lead to doing nothing is better than nonsence.

Question - 2

How do we measure the cost of preferred stock and cost of common stock equity of a firm?

Preferred stock: These stocks had an unique feature of promised dividend. So depending on market price of the share, we can easily calculate the expected required rate for preferred stock. For example a preferred stock with 100 face value had 10% dividend as coupon rate and is currently selling at 120 per share. In such case

Kp = D/ SV * 100 = 10 / 120 * 100 = 8.33%

Common stock : These are also called variable dividend securities, because dividend is not predetermined for them. In such case valuation or determination of expectations is quite difficult. Hence several models were developed over the time for computing the cost of common stock. Best of such methods is CAPM method

Ke = Risk free rate + Beta ( Market risk premium)